OpenStudy (anonymous):

Give the direction of -6i+5j to the nearest degree

OpenStudy (anonymous):

They want you to convert the vector to an angle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve for the angle.

OpenStudy (anonymous):

Remember to use degree mode.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

just plug in tan 5/-6 into the calc?

OpenStudy (anonymous):

you also need to do the inverse tangent.

OpenStudy (anonymous):

I get -.8333333333

OpenStudy (anonymous):

You have to solve for the angle, so you need to do the inverse tangent.

OpenStudy (anonymous):

\[ \tan\theta =\frac{5}{-6} \implies \theta = \tan^{-1}\left(\frac{5}{-6}\right) \]

OpenStudy (unklerhaukus):

dont forget to add 180°

OpenStudy (anonymous):

got it! thanks!

OpenStudy (anonymous):

will it always be tangent?

OpenStudy (anonymous):

Remember SOH CAH TOA, but when it comes to vectors, you're going to use inverse tangent cause you're only dealing with TOA.

OpenStudy (unklerhaukus):

well the unit vectors will usually be perpendicular so they will always be the adjacent and opposite sides (never hypotenuse )

OpenStudy (anonymous):

ok got it!